# Expansion Span: Combining Fading Memory and Retrieval in Hybrid State Space Models

### v1

@@ 7.1 Fine-Tuning on Natural Language + Code
Next, we consider fine-tuning on a dataset that augments natural language with a code dataset. In particular, we construct a dataset consisting of 70% natural language, and 30% C++ code extracted from the Lots of Code [39] dataset. We provide PG-19 validation PPL results, as well as results on tasks from the LM Evaluation Harness suite in Tab. 5. Similar to fine-tuning on natural language only (as in Tab. 2), we observe that fine-tuning with SE-Attn yields the strongest performance on downstream tasks. We provide performance on RULER in Figure 9. We again observe that fine-tuning with SE-Attn yields the strongest perfromance compared to other efficient Attention layers. Moreover, compared to fine-tuning only on natural language, here we observe a greater improvement on the Variable Tracking task. As explained in Sec. 11.4, this task requires the model to keep track of the values of variables that are defined and overridden throughout the input context, and must then return the variables equal to some value. This requires strong recall capabilities, which fine-tuning with SE-Attn enables, and is amplified by the use of code data in the fine-tuning data mix.
@@
7.2 Fine-Tuning on PG-19
Next, we consider fine-tuning on the PG-19 [27] dataset. We provide perplexity results on the validation set, along with LM Harness task results in Tab. 6. Compared to fine-tuning on a different natural language dataset, and a natural language + code as in Tab. 2 and Tab. 5, here we obtain lower perplexity scores due to the lack of distribution shift. Interestingly, compared to fine-tuning on the previously mentioned datasets, we observe a slight degradation on LM Harness tasks. Moreover, in Figure 10 we plot RULER performance when fine-tuning on PG-19, and here we also see a decay in performance compared to fine-tuning on the previous datasets. This suggests that the PG-19 dataset may be too far out of distribution for these long-context retrieval task. Nevertheless, we see that fine-tuning with SE-Attn yields the best performance across all of these tasks.
@@ 7.2 Fine-Tuning on PG-19
Figure 10: Mamba-2-Hybrid RULER benchmark fine-tuned on PG-19. We fine-tune Mamba-2-Hybrid with different Attention layers on the PG-19 [27] dataset and then evaluate on RULER. Compared to fine-tuning on other natural language datasets with a greater variety of text as in Figure 8, here we observe a degradation in performance across all models, likely due to a distribution shift in the PG-19 data and the RULER tasks.
7.3 Evaluating with Efficient Attention
@@
In Sec. 4.2, we explained how we fine-tune our models using an efficient Attention mechanism, and then evaluate with full attention. In Tab. 1, we showed that evaluating our models with efficient Attention mechanisms produced strong results on perplexity benchmarks. However, we did not observe similarly positive results on more complex tasks, such as those in RULER. As illustrated in Figure 11, using the same efficient Attention layer used during fine-tuning does not perform as well as evaluating with Full-Attn (we omit results for models fine-tuned with $S^{2}$-Attn and evaluated with $S^{2}$-Attn as the implementation of $S^{2}$-Attn does not support evaluating with it at arbitrary sequence lengths). Hence, our training pipeline involves training a model with an efficient Attention layer, and reverting to Full-Attn during inference. Moreover, we note that perplexity is again a misleading metric in regard to measuring a model’s performance on long context tasks. Though we observed strong performance on perplexity when evaluating with these efficient Attention layers, this did not translate to stronger performance on tasks that require a strong recall capability, such as those in RULER.
Figure 11: Evaluating with efficient Attention mechanisms on RULER does not do as well as evaluating with standard full Attention. We fine-tune Mamba-2-Hybrid with SW-Attn (a) and SE-Attn (b) and then evaluate on the NIAH-Single-1 RULER task using either the same Attention layer used for fine-tuning, or Full-Attn. We observe that fine-tuning with efficient Attention layers (SE-Attn and SW-Attn) and then using Full-Attn during evaluation yields better results. Lines with a circle marker denote models fine-tuned with an efficient Attention mechanism, and then evaluated with the same Attention mechanism; lines with a diamond marker were fine-tuned with an efficient Attention mechanism, but evaluated with Full-Attn. We omit results for evaluating models fine-tuned with $S^{2}$-Attn as this Attention mechanism does not support arbitrary sequence lengths during inference.
8 Retrieval with Landmark Tokens
@@
Landmark Attention [23] was recently introduced as way for Transformer models to process long sequences by inserting ‘landmark’ tokens into the sequence whose representations would then be used as summaries of the blocks of tokens that came before them. At a high level, our approach in SE-Attn is similar. However, we simplify the process of compressing blocks of tokens by forgoing the use of landmark tokens and instead using Attention to summarize them, as described in Sec. 3.3. Moreover, to learn which blocks to retrieve, we do not rely on a complex Grouped Softmax function, and instead use a simple Cross-Attention score to ascertain relevance. In this way, we implement retrieval natively into the model’s architecture.
@@
We consider a variant of SE-Attn, which we refer to as SE-Attn-LM, that is inspired by Landmark Attention. Namely, instead of using our Attention-based compression to construct summaries of memory blocks, we insert a non-learnable ‘landmark’ token into each memory block, and use the cross-attention between this token and the memory tokens as the summary of the memory block. We compare the performance of this variant to our summaries computing using average pooling of Attention scores (see Equation 7) in Figure 12. Here, we see that SE-Attn yields better performance. We suspect this is because using a non-learnable landmark token to summarize memory blocks is too challenging of a task to accomplish using standard LoRA. While full fine-tuning (without LoRA) may improve the performance of SE-Attn-LM, this is beyond the scope of our work as we prioritize efficiency.
@@
Figure 12: Summarizing memory blocks via average pooling of attention yields stronger performance than summarizing them using ‘landmark’ tokens. We fine-tune Mamba-2-Hybrid using our SE-Attn, and SE-Attn-LM, which summarizes memory blocks with a landmark token (similar to [23]) instead of using the average of the Self Attention output of memory blocks, as in SE-Attn. We find that our simpler SE-Attn produces a stronger model, likely due to the easier training task, which does not require adapting the model to leverage landmark tokens for compression.
9 Training Details
@@
Our fine-tuning recipe largely follows [3], with the exception of using a larger learning rate for SE-Attn, SW-Attn, and Full-Attn models ($2\times 10^{-4}$ vs. the default $2\times 10^{-5}$ for $S^{2}$). We found that using a larger learning rate for $S^{2}$ did not improve its performance, as shown in Figure 13, so we used the default $2\times 10^{-5}$ learning rate for all $S^{2}$ fine-tuning experiments.
@@
All of our experiments are conducted on a single 8xA100 node. We fine-tune for 1000 steps with a total of 0.5B tokens. We fine-tune Mamba-2-H with 8192 tokens with 8 accumulation steps and a per-device batch size of 1. We fine-tune Llama1 with 16384 tokens with 4 accumulation steps and a per-device batch size of 1. We use FlashAttention-2 [5] and DeepSpeed Stage 2 [28].
@@
Figure 13: A larger learning rate does not improve the performance of $S^{2}$. Here we fine-tune a Mamba-2-Hybrid model using $S^{2}$-Attn with two different learning rates: $2\times 10^{-4}$ and $2\times 10^{-5}$. The learning rate used in [3] is $2\times 10^{-5}$, which we found to work well. For all other Attention layers, we found $2\times 10^{-4}$ to offer a slight improvement over $2\times 10^{-5}$.
10 Runtime Analysis
@@
The runtime of SE-Attn is lower than Full-Attn and $S^{2}$-Attn on large contexts. For an input with sequence length $L$, SE-Attn first constructs $U=\frac{L}{S}$ memory blocks of size $S$. Attention is applied on each. Thus, the complexity of this operation is $O(d_{\textrm{model}}US^{2})=O(d_{\textrm{model}}LS)$. Next, the input sequence is split into $T=\frac{L}{M}$ chunks of size $M$. Cross-Attention is then applied between each chunk’s query and each memory block’s compressed representation; the cost of this is $O(d_{\textrm{model}}MTU)=O(d_{\textrm{model}}\frac{L^{2}}{S})$. Cross-Attention is then applied between each chunk’s query tokens and the concatenations of the chunk’s key tokens and memory tokens. Since SE-Attn retrieves $K$ blocks with $S$ tokens in each, the cost of this Cross-Attention for each chunk is $O(d_{\textrm{model}}M(SK+M))=O(d_{\textrm{model}}M^{2})$ since $SK<M$. The cost for all chunks is therefore $O(d_{\textrm{model}}TM^{2})=O(d_{\textrm{model}}LM)$. The total cost of SE-Attn is therefore $O(d_{\textrm{model}}LS+d_{\textrm{model}}LM+d_{\textrm{model}}\frac{L^{2}}{S})$. For sufficiently large $S$, the runtime of SE-Attn is faster than Full-Attn and $S^{2}$-Attn, especially on large contexts, as shown empirically in Sec. 4.7.
11 RULER Task Definitions
@@
The RULER benchmark [16] consists of four different task categories: retrieval, multi-hop tracing, aggregation, and question answering. In this paper, we focus only on the retrieval, multi-hop tracing, and aggregation tasks (we provide question answering results on the LM Evaluation Harness benchmark). These three categories span eleven different tasks as explained below.
11.1 Needle-in-a-Haystack (NIAH) Tasks
@@
  • type_haystack (TH): This specifies the type of noise to embed the key in. The choices are “repeat” which constructs noise as in [23], “essay” which will use sentences from the Paul Graham essays [20], or “needle” in which case each sentence will define a new key-value pair.
@@
Table 7: RULER NIAH definitions. The ‘Needle-in-a-Haystack’ (NIAH) tasks in the RULER benchmarks are defined by 6 parameters which modulate the difficulty of the tasks. We consider 8 different NIAH tasks as defined above (these are the default NIAH tasks in the RULER library).
11.2 Multi-hop Tracing Tasks
RULER considers a “variable tracking” task that is a form of coreference resolution. In this task, a sequence of variables are defined throughout noisy text as in [23]. New variables are defined as previous ones, and a final value is assigned to a particular variable. The goal is to be able to trace back which variables have also been assigned the final value, i.e., determine which variables refer to the final value. We use the default num_chains=1 and num_hops=4 parameters
11.3 Aggregation Tasks
@@
11.4 Aggregating RULER Tasks

### v2

@@ E.1 Fine-Tuning on Natural Language + Code
Next, we consider fine-tuning on a dataset that augments natural language with a code dataset. In particular, we construct a dataset consisting of 70% natural language, and 30% C++ code extracted from the Lots of Code [Zavadskyy(2017)] dataset. We provide PG-19 validation PPL results, as well as results on tasks from the LM Evaluation Harness suite in Table 6. Similar to fine-tuning on natural language only (as in Table 1), we observe that fine-tuning with SE-Attn yields the strongest performance on downstream tasks. We provide performance on RULER in Figure 7. We again observe that fine-tuning with SE-Attn yields the strongest perfromance compared to other efficient Attention layers. Moreover, compared to fine-tuning only on natural language, here we observe a greater improvement on the Variable Tracking task. As explained in Section I.4, this task requires the model to keep track of the values of variables that are defined and overridden throughout the input context, and must then return the variables equal to some value. This requires strong recall capabilities, which fine-tuning with SE-Attn enables, and is amplified by the use of code data in the fine-tuning data mix.
@@
E.2 Fine-Tuning on PG-19
Next, we consider fine-tuning on the PG-19 [Rae et al.(2019b)Rae, Potapenko, Jayakumar, and Lillicrap] dataset. We provide perplexity results on the validation set, along with LM Harness task results in Table 7. Compared to fine-tuning on a different natural language dataset, and a natural language + code dataset as in Table 1 and Table 6, here we obtain lower perplexity scores due to the lack of distribution shift. Interestingly, compared to fine-tuning on the previously mentioned datasets, we observe a slight degradation on LM Harness tasks. Moreover, in Figure 8 we plot RULER performance when fine-tuning on PG-19, and here we also see a decay in performance compared to fine-tuning on the previous datasets. This suggests that the PG-19 dataset may be too far out of distribution for these long-context retrieval task. Nevertheless, we see that fine-tuning with SE-Attn yields the best performance across all of these tasks.
@@ E.2 Fine-Tuning on PG-19
Figure 8: Mamba-2-Hybrid RULER benchmark fine-tuned on PG-19. We fine-tune Mamba-2-Hybrid with different Attention layers on the PG-19 [Rae et al.(2019b)Rae, Potapenko, Jayakumar, and Lillicrap] dataset and then evaluate on RULER. Compared to fine-tuning on other natural language datasets with a greater variety of text as in Figure 6, here we observe a degradation in performance across all models, likely due to a distribution shift in the PG-19 data and the RULER tasks.
E.3 Additional Long-context Benchmarks
In this section, we evaluate our Mamba-2-Hybrid model fine-tuned on only language data on additional long-context tasks. First, we test our models’ long-context understanding on the LongBench benchmark and report results in Table 8. Here, we observe that our fine-tuned models improve over the pre-trained baseline models, suggesting that these models’ long-context understanding has improved. Moreover, we see that models fine-tuned with our SE-Attn match the performance of the model fine-tuned with Full-Attn .
We previously considered RULER as a means of assessing our models’ in-context recall capabilities. While RULER can assess our models’ performance on synthetic in-context recall tasks, we also want to evaluate our their performance on “real-world” in-context tasks. As such, in Table 9, we evaluate on the tasks considered in [Arora et al.(2024)Arora, Timalsina, Singhal, Spector, Eyuboglu, Zhao, Rao, Rudra, and Ré]. We observe that fine-tuning with our SE-Attn is able to match the performance of fine-tuning with Full-Attn on these tasks.
Attention Single-Doc QA ($\uparrow$) Multi-Doc QA ($\uparrow$) Summarization ($\uparrow$) Few-shot ($\uparrow$) Code ($\uparrow$) NQA QQA MFQ HQA 2WM Mus GvR QMS MNs TRC TQA LCC RBP Avg. Non-fine-tuned 0.92 3.78 13.92 4.72 8.26 1.54 9.65 3.71 23.23 59.0 63.22 60.55 31.99 21.88 Full-Attn 3.25 7.15 16.35 8.03 9.51 3.69 25.4 17.64 26.05 64.5 82.09 56.67 55.58 28.92 SW-Attn 2.29 7.52 17.07 7.72 9.4 3.41 18.44 11.87 23.94 64.5 81.48 58.33 53.06 27.62 $S^{2}$-Attn 1.39 5.58 16.43 5.31 7.88 1.83 13.64 5.64 26.86 57.0 72.97 56.62 36.87 23.69 SE-Attn 3.19 9.37 16.0 8.31 9.31 3.94 24.65 14.39 26.86 66.5 80.82 58.71 53.65 28.90
Table 8: SE-Attn matches the performance of Full-Attn on long-context understanding tasks. We fine-tune Mamba-2-Hybrid models using various attention mechanisms on 14 LongBench [Bai et al.(2024)Bai, Lv, Zhang, Lyu, Tang, Huang, Du, Liu, Zeng, Hou, Dong, Tang, and Li] tasks.
Attention SWDE SQUADv2 FDA TriviaQA NQ Drop Avg
Non-fine-tuned 85.60 50.11 76.41 22.60 6.62 3.20 40.76
Full-Attn 85.24 50.09 80.22 25.75 6.79 3.55 41.94
SW-Attn 84.61 50.10 78.31 25.51 6.76 3.62 41.48
$S^{2}$-Attn 86.41 50.11 81.49 23.63 6.84 3.32 41.97
SE-Attn 85.96 50.09 82.12 26.39 7.17 3.76 42.58
Table 9: SE-Attn matches the performance of Full-Attn on real-world in-context tasks. We fine-tune Mamba-2-Hybrid models with different attention layers. We do not truncate inputs to 2K tokens when evaluating as was done in [Arora et al.(2024)Arora, Timalsina, Singhal, Spector, Eyuboglu, Zhao, Rao, Rudra, and Ré].
E.4 Evaluating with Efficient Attention
@@
In Appendix A, we explained how we fine-tune our models using an efficient Attention mechanism, and then evaluate with full attention. In Table 2, we showed that evaluating our models with efficient Attention mechanisms produced strong results on perplexity benchmarks. However, we did not observe similarly positive results on more complex tasks, such as those in RULER. As illustrated in LABEL:fig:ft_attn_niah, using the same efficient Attention layer used during fine-tuning does not perform as well as evaluating with Full-Attn (we omit results for models fine-tuned with $S^{2}$-Attn and evaluated with $S^{2}$-Attn as the implementation of $S^{2}$-Attn does not support evaluating with it at arbitrary sequence lengths). Hence, our training pipeline involves training a model with an efficient Attention layer, and reverting to Full-Attn during inference. Moreover, we again note that perplexity is a misleading metric in regard to measuring a model’s performance on long context tasks. Though we observed strong performance on perplexity when evaluating with these efficient Attention layers, this did not translate to stronger performance on tasks that require a strong recall capability, such as those in RULER.
Appendix F Retrieval with Landmark Tokens
@@
Landmark Attention [Mohtashami and Jaggi(2023)] was recently introduced as a way for Transformer models to process long sequences by inserting “landmark” tokens into the sequence whose representations would then be used as summaries of the blocks of tokens that came before them. At a high level, our approach in SE-Attn is similar. However, we simplify the process of compressing blocks of tokens by forgoing the use of landmark tokens and instead using Attention to summarize them, as described in Section 3.3. Moreover, to learn which blocks to retrieve, we do not rely on a complex Grouped Softmax function, and instead use a simple Cross-Attention score to ascertain relevance. In this way, we implement retrieval natively into the model’s architecture.
@@
In this section, we consider a variant of SE-Attn, which we refer to as SE-Attn-LM, that is inspired by Landmark Attention. Namely, instead of using our Attention-based compression to construct summaries of memory blocks, we insert a non-learnable “landmark” token into each memory block, and use the cross-attention between this token and the memory tokens as the summary of the memory block. We compare the performance of this variant to our summaries computed using average pooling of Attention scores (see Section 3.3) in Figure 10. Here, we see that SE-Attn yields better performance. We suspect this is because using a non-learnable landmark token to summarize memory blocks is too challenging of a task to accomplish using standard LoRA. While full fine-tuning (without LoRA) may improve the performance of SE-Attn-LM, this is beyond the scope of our work as we prioritize efficiency.
@@
Figure 10: Summarizing memory blocks via average pooling of attention yields stronger performance than summarizing them using “landmark” tokens. We fine-tune Mamba-2-Hybrid using our SE-Attn, and SE-Attn-LM, which summarizes memory blocks with a landmark token (similar to [Mohtashami and Jaggi(2023)]) instead of using the average of the Self Attention output of memory blocks, as in SE-Attn. We find that our simpler SE-Attn produces a stronger model, likely due to the easier training task, which does not require adapting the model to leverage landmark tokens for compression.
Appendix G Training Details
@@
Our fine-tuning recipe largely follows [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia], with the exception of using a larger learning rate for SE-Attn, SW-Attn, and Full-Attn models ($2\times 10^{-4}$ vs. the default $2\times 10^{-5}$ for $S^{2}$). We found that using a larger learning rate for $S^{2}$ did not improve its performance, as shown in Figure 11, so we used the default $2\times 10^{-5}$ learning rate for all $S^{2}$ fine-tuning experiments.
@@
All of our experiments are conducted on a single 8xA100 node. We fine-tune for 1000 steps on a total of 0.5B tokens. We fine-tune Mamba-2-H with a context size of 8192 with 8 accumulation steps and a per-device batch size of 1. We fine-tune Llama1 with with a context size of 16384 tokens with 4 accumulation steps and a per-device batch size of 1. We use FlashAttention-2 [Dao(2024)] and DeepSpeed Stage 2 [Rasley et al.(2020)Rasley, Rajbhandari, Ruwase, and He].
@@
Figure 11: A larger learning rate does not improve the performance of $S^{2}$. Here we fine-tune a Mamba-2-Hybrid model using $S^{2}$-Attn with two different learning rates: $2\times 10^{-4}$ and $2\times 10^{-5}$. The learning rate used in [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia] is $2\times 10^{-5}$, which we found to work well. For all other Attention layers, we found $2\times 10^{-4}$ to offer a slight improvement over $2\times 10^{-5}$.
Appendix H Runtime and Memory Analysis
@@
The runtime of SE-Attn is lower than Full-Attn and $S^{2}$-Attn on large contexts. For an input with sequence length $L$, SE-Attn first constructs $U=\frac{L}{S}$ memory blocks of size $S$. Attention is applied on each. Thus, the complexity of this operation is $O(d_{\textrm{model}}US^{2})=O(d_{\textrm{model}}LS)$. Next, the input sequence is split into $T=\frac{L}{M}$ chunks of size $M$. Cross-Attention is then applied between each chunk’s query and each memory block’s compressed representation; the cost of this is $O(d_{\textrm{model}}MTU)=O(d_{\textrm{model}}\frac{L^{2}}{S})$. Cross-Attention is then applied between each chunk’s query tokens and the concatenations of the chunk’s key tokens and memory tokens. Since SE-Attn retrieves $K$ blocks with $S$ tokens in each, the cost of this Cross-Attention for each chunk is $O(d_{\textrm{model}}M(SK+M))=O(d_{\textrm{model}}M^{2})$ since $SK<M$. The cost for all chunks is therefore $O(d_{\textrm{model}}TM^{2})=O(d_{\textrm{model}}LM)$. The total cost of SE-Attn is therefore $O(d_{\textrm{model}}LS+d_{\textrm{model}}LM+d_{\textrm{model}}\frac{L^{2}}{S})$. For sufficiently large $S$, the runtime of SE-Attn is faster than Full-Attn and $S^{2}$-Attn, especially on large contexts, and is similar to that SW-Attn.
Appendix I RULER Task Definitions
@@
The RULER benchmark [Hsieh et al.(2024)Hsieh, Sun, Kriman, Acharya, Rekesh, Jia, Zhang, and Ginsburg] consists of four different task categories: retrieval, multi-hop tracing, aggregation, and question answering. In this paper, we focus only on the retrieval, multi-hop tracing, and aggregation tasks (we provide question answering results on the LM Evaluation Harness benchmark). These three categories span eleven different tasks as explained below.
I.1 Needle-in-a-Haystack (NIAH) Tasks
@@
  • type_haystack (TH): This specifies the type of noise to embed the key in. The choices are “repeat” which constructs noise as in [Mohtashami and Jaggi(2023)], “essay” which will use sentences from the Paul Graham essays [Kamradt(2023)], or “needle” in which case each sentence will define a new key-value pair.
@@
Table 10: RULER NIAH definitions. The “Needle-in-a-Haystack” (NIAH) tasks in the RULER benchmarks are defined by 6 parameters which modulate the difficulty of the tasks. We consider 8 different NIAH tasks as defined above (these are the default NIAH tasks in the RULER library).
I.2 Multi-hop Tracing Tasks
RULER considers a “variable tracking” task that is a form of coreference resolution. In this task, a sequence of variables are defined throughout noisy text as in [Mohtashami and Jaggi(2023)]. New variables are defined as previous ones, and a final value is assigned to a particular variable. The goal is to be able to trace back which variables have also been assigned the final value, i.e., determine which variables refer to the final value. We use the default num_chains=1 and num_hops=4 parameters
I.3 Aggregation Tasks
@@
I.4 Aggregating RULER Tasks
@@
Appendix J Task Abbreviations
LM Harness: Hella=HellaSwag, LAMB=LAMBADA, WG=WinoGrande, SQA=ScrollsQAsper, SNQA=ScorllsNarrativeQA.
LongBench: NQA=NarrativeQA, QQA=QasperQA, MFQ=MultiFieldQA, HQA=HotpotQA,
2WM=2WikiMultihopQA, Mus=Musiqe, GvR=GovReport, QMS=QMSum, MNs=MultiNews, TRC=TREC, TQA=TriviaQA, SSM=SamSum, RBP=RepoBench-P.